A capacitor has a charge of 9 microcoulomb when connected with a 12 V battery. Determine the capacitance.


1
Expert's answer
2022-03-13T18:49:29-0400

The capacitance is given as follows:



"C = \dfrac{Q}{V}"

where "Q = 9\times 10^{-9}C, V = 12V". Thus, obtain:



"C = \dfrac{9\times 10^{-9}C}{12V} = 7.5\times 10^{-10}F"

Answer. "7.5\times 10^{-10}F".